OpenStudy (anonymous):

Find the equation of the line that goes through (0,-4) and has a slope of -6

OpenStudy (gitface):

Use the equation; \[y - y _{1} = m(x-x _{1})\] Where y1 is the given y-coordinate, x1 is the given x-coordinate, m is the slope \[y - (-4) = -6(x-0)\] Solve for y.

OpenStudy (anonymous):

using point-slope form y = -6(x-0) -4

OpenStudy (anonymous):

awesome, thank you! i will try it

OpenStudy (anonymous):

so would the answer be y=-6x-4?

OpenStudy (anonymous):

so would the answer be y=-6x-4?

OpenStudy (gitface):

That is correct.

OpenStudy (anonymous):

thank youuuU!!!! awesome